OpenStudy (anonymous):

Write a system of equations for the following situation. A 5,000-seat theater has tickets for sale at $25 and $40. How many tickets should be sold at each price for a sellout performance to generate a total revenue of $152,000?

OpenStudy (anonymous):

Let \(x\) be the number of $25 tickets needing to be sold and let \(y\) be the same for the $40 tickets. $$25x + 40y = 152000$$ This is the first equation, and is evident from total revenue.

OpenStudy (anonymous):

so the 2nd would would be 25x+40y=5000 ?

OpenStudy (ganpat):

naaa.. 2nd equation would be x + y = 5000